For the spring-mass system shown below with the mass sliding on a frictionless floor, A = 1.0 m, the spring constant k = 2.0 N/m, and the mass m = 2.0 kg. The period of oscillation T isQUESTION 18 For the spring-mass system shown below with the mass sliding on a frictionless floor, A = 1.0 m, the spring const
Answer #1

mass of spring (m) = 2 kg K = 2 Nim Т. АТС 2T/W - w= 2 w=1 rad/SEL T= 20 I - 212 T= 2x3.14 T = 6-28 T= 6-3 sec a Answor